The rate equation for a reaction: A → B is r=K[A]0 . If the initial concentration of the reactant is a moldm-3 , the half-life period of the reaction is?


1
Expert's answer
2021-09-07T14:39:44-0400

Given:

The rate equation for the reaction is r=k[A]0.

Initial concentration of the reactant = a moldm-3


The rate law of the reaction is expressed as:

Rate=k[A]x

Here, x is the order of the reaction and k is the rate constant for the reaction.

According to the given rate law, r=k[A]0, the order of the reaction is zero order.


For a zero order reaction, the rate constant is given by:

"k=\frac{[A]o−[A]}{t}......(1)"

Here, [A]o is the initial concentration of the reactant and [A]

A is the final concentration of the reactant after time t.


Half-life period is the time period at which the final concentration of reactant becomes half of the initial concentration of the reactant. Hence,"[A]=\frac{1}{2}[A]_o."


Substitute the given values in equation (1).

"k=\frac{[A]_o−\frac{[A]_o}{2}}{t_{\frac{1}{2}}}\\k=\frac{2[A]_o−[A]_o}{2t_{\frac{1}{2}}}\\t_{\frac{1}{2}}=\frac{[A]_o}{2k}\\t_{\frac{1}{2}}=\frac{a}{2k}"

Therefore, the half-life period of the reaction is "\frac{a}{

2k}."